# Retraining with Predicted Hard Labels Provably Increases Model Accuracy

### v1

@@ 4 Full Retraining in the Presence of Label Noise: Theoretical Analysis
We are given $n>d$ labeled examples $\mathcal{X}_{L}:=\{{\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{n}\}$ and $n$ unlabeled examples $\mathcal{X}_{U}:=\{{\mathbf{x}}^{\prime}_{1},\ldots,{\mathbf{x}}^{\prime}_{n}\}$ drawn i.i.d. from $\mathcal{D}$. For $j\in[n]$, let $\mathrm{y}_{j}=\text{sign}\big{(}\langle{\mathbf{x}}_{j},\bm{\theta}^{\ast}\rangle\big{)}$ be the ground truth label of ${\mathbf{x}}_{j}$. We have access to a noisy label source which gives us a randomly corrupted version of $\mathrm{y}_{j}$, namely $\widehat{\mathrm{y}}_{j}$, instead of $\mathrm{y}_{j}$. Specifically, for some $p<\frac{1}{2}$, we have: [^5] In the context of $\epsilon$-labelDP with randomized response, $p=\frac{1}{1+e^{\epsilon}}$.
$\widehat{\mathrm{y}}_{j}=\begin{cases}\mathrm{y}_{j}\text{ w.p. }1-p\\ -\mathrm{y}_{j}\text{ w.p. }p.\end{cases}$ (1)
Let $\mathcal{T}:=\{({\mathbf{x}}_{j},\widehat{\mathrm{y}}_{j})\}_{j\in[n]}$ denote the labeled part of the training data. We use $\mathcal{T}$ for vanilla training. Then, we label $\mathcal{X}_{U}$ using the classifier learned with vanilla training and use this for retraining. Ideally, we would have liked to analyze retraining over $\mathcal{X}_{L}$ and not have $\mathcal{X}_{U}$ at all, but this is complicated and tedious due to independence not holding. In order to circumvent this extra layer of complication in the analysis with $\mathcal{X}_{L}$, we analyze retraining over $\mathcal{X}_{U}$ (wherein independence obviates the said complication). We believe that the main insights should not change much.
4.1 Vanilla Training
Here we train with the given noisy labels as is. Our objective function with the squared loss is $\frac{1}{2n}\sum_{j=1}^{n}\big{(}\widehat{\mathrm{y}}_{j}-\bm{\theta}^{\top}{\mathbf{x}}_{j}\big{)}^{2}$ and its minimizer is $\bar{\bm{\theta}}_{0}=\Big{(}\frac{1}{n}\sum_{k=1}^{n}{\mathbf{x}}_{k}{\mathbf{x}}_{k}^{\top}\Big{)}^{-1}\Big{(}\frac{1}{n}\sum_{j=1}^{n}\widehat{\mathrm{y}}_{j}{\mathbf{x}}_{j}\Big{)}$. But instead of considering $\bar{\bm{\theta}}_{0}$ as our learned separator, we will focus on
$\widehat{\bm{\theta}}_{0}=\frac{1}{\sigma^{2}}\Bigg{(}\frac{1}{n}\sum_{j=1}^{n}\widehat{\mathrm{y}}_{j}{\mathbf{x}}_{j}\Bigg{)}.$ (2)
Observe that the empirical covariance matrix $\frac{1}{n}\sum_{k=1}^{n}{\mathbf{x}}_{k}{\mathbf{x}}_{k}^{\top}$ in $\bar{\bm{\theta}}_{0}$ is replaced by its expected value, viz., $\sigma^{2}{\mathbf{I}}_{d}$ in $\widehat{\bm{\theta}}_{0}$. The reason for doing this is that the analysis with $\bar{\bm{\theta}}_{0}$ eventually reduces down to analyzing with $\widehat{\bm{\theta}}_{0}$ by performing some tedious math first (due to the empirical covariance matrix) which does not yield any meaningful insights. For completeness, we discuss how the analysis can be done with $\bar{\bm{\theta}}_{0}$ in Appendix G. Our predicted label for a sample ${\mathbf{x}}$ with $\widehat{\bm{\theta}}_{0}$ is $\textup{sign}\big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle\big{)}$ (instead of just $\langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle$ which would have been the case in a regression problem).
Theorem 1 (Vanilla Training: Probability of Predicted Label = True Label).
Consider a sample ${\mathbf{x}}\notin\mathcal{X}_{L}$. Let $\beta({\mathbf{x}})$ denote the angle between ${\mathbf{x}}$ and $\bm{\theta}^{\ast}$. Then our predicted label for ${\mathbf{x}}$, viz., $\textup{sign}\big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle\big{)}$ is equal to the ground truth label of ${\mathbf{x}}$, viz., $\textup{sign}\big{(}\langle{\mathbf{x}},\bm{\theta}^{\ast}\rangle\big{)}$ with a probability (over the randomness of $\mathcal{T}$) of at least $\alpha({\mathbf{x}}):=1-\exp\Big{(}-\frac{n\mu^{2}}{4}(1-2p)^{2}\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Big{)}$ and at most $\widehat{\alpha}({\mathbf{x}}):=1-\frac{1}{3\sqrt{2\pi}}{\exp\Big{(}-{\big{(}1+\sqrt{n}\mu(1-2p)\big{)}^{2}\textup{cot}^{2}\big{(}\beta({\mathbf{x}})\big{)}}\Big{)}}$.
The proof of Theorem 1 is in Appendix B. Observe that $\alpha({\mathbf{x}})$ and $\widehat{\alpha}({\mathbf{x}})$ have essentially the same dependence on $n$, $p$ and $\mu$; so, this dependence is tight. Also, notice that the learned classifier $\widehat{\bm{\theta}}_{0}$ is more likely to be wrong on the samples that are less aligned with (or closer to orthogonal to) the ground truth separator, i.e., $\bm{\theta}^{\ast}$. We can view $\widehat{\bm{\theta}}_{0}$ as a noisy label provider (a.k.a. pseudo-labeler) where the degree of label noise is non-uniform or sample-dependent unlike the original noisy source used to learn $\widehat{\bm{\theta}}_{0}$. Specifically, for a sample ${\mathbf{x}}$ with true label $\mathrm{y}=\text{sign}\big{(}\langle{\mathbf{x}},\bm{\theta}^{\ast}\rangle\big{)}$ and predicted label $\widetilde{\mathrm{y}}=\textup{sign}\big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle\big{)}$, we have:
[…]
We will now provide lower and upper bounds on the population accuracy of the classifier $\widehat{\bm{\theta}}_{0}$; we denote this by $\text{acc}(\widehat{\bm{\theta}}_{0})$ and it is defined as:
$\text{acc}(\widehat{\bm{\theta}}_{0}):=\mathbb{P}_{{\mathbf{x}}\sim\mathcal{D},\mathcal{T}}\Big{(}\textup{sign}\big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle\big{)}=\textup{sign}\big{(}\langle{\mathbf{x}},\bm{\theta}^{\ast}\rangle\big{)}\Big{)}.$
Theorem 2 (Vanilla Training: Population Accuracy).
We have:
$\textup{acc}(\widehat{\bm{\theta}}_{0})\geq 1-\exp\big{(}-\Theta(d)\big{)}-\exp\Bigg{(}-\frac{n{\mu^{2}}\gamma^{2}(1-2p)^{2}}{\Theta(d)}\Bigg{)},$
and
$\textup{acc}(\widehat{\bm{\theta}}_{0})\leq 1-\Theta(1)\exp\Bigg{(}-\frac{n\mu^{4}(1-2p)^{2}}{\Theta(d)}\Bigg{)}.$
The proof of Theorem 2 is in Appendix C.
Remark 1 (Tightness of Accuracy Bounds).
@@ 4.1 Vanilla Training
Suppose $\gamma=\Theta(\mu)$. When $n\leq\Theta\Big{(}\frac{d^{2}}{(1-2p)^{2}\mu^{4}}\Big{)}$, the lower and upper bounds for $\textup{acc}(\widehat{\bm{\theta}}_{0})$ in Theorem 2 essentially match (modulo constant factors).
Based on Theorem 2, we have the following corollary.
Corollary 1 (Vanilla Training: Sample Complexity).
@@
For any $\delta>(1+\Theta(1))\exp(-\Theta({d}))$, $n=\Omega\Big{(}\frac{\log{1}/{\delta}}{(1-2p)^{2}}\big{(}\frac{d}{\mu^{2}\gamma^{2}}\big{)}\Big{)}$ ensures that $\textup{acc}(\widehat{\bm{\theta}}_{0})>1-\delta$. Specifically, when $p>(1+\Theta(1))\exp(-\Theta({d}))$, $n=\Omega\Big{(}\frac{\log{1}/{p}}{(1-2p)^{2}}\big{(}\frac{d}{\mu^{2}\gamma^{2}}\big{)}\Big{)}$ ensures that $\textup{acc}(\widehat{\bm{\theta}}_{0})>1-p$, i.e., our learned classifier $\widehat{\bm{\theta}}_{0}$ has better accuracy than the source providing noisy labels (used to learn $\widehat{\bm{\theta}}_{0}$).
We will now present an information-theoretic lower bound on the sample complexity of any classifier to argue the optimality of the result in Corollary 1 with respect to the dependence on $d$ and $p$.
Theorem 3 (Information-Theoretic Lower Bound on Sample Complexity).
With a slight generalization of notation, let $\textup{acc}(\widehat{\bm{\theta}};\bm{\theta}^{*})$ denote the accuracy of the classifier $\widehat{\bm{\theta}}$, when the ground truth model is $\bm{\theta}^{*}$. For any classifier $\widehat{\bm{\theta}}$ learned from $\mathcal{T}:=\{({\mathbf{x}}_{j},\widehat{\mathrm{y}}_{j})\}_{j\in[n]}$, in order to achieve $\inf_{\bm{\theta}^{\ast}\in S^{d-1}}\textup{acc}(\widehat{\bm{\theta}};\bm{\theta}^{*})\geq 1-\delta$, where $\delta<\frac{1}{5}$, $n=\Omega\left(\frac{(1-5\delta)}{(1-2p)^{2}}d\right)$ is necessary in our problem setting.
It is worth mentioning that there is a similar lower bound in Gentile and Helmbold, (1998) for a different classification setting. In contrast, Theorem 3 is tailored to our setting and moreover, the proof technique is also different. Specifically, for the proof of Theorem 3, we follow a standard technique in proving minimax lower bounds which is to reduce the problem of interest to an appropriate multi-way hypothesis testing problem; this is accompanied by the application of the conditional version of Fano’s inequality and some ideas from high-dimensional geometry. This proof is in Appendix D.
Remark 2 (Optimality of Sample Complexity).
Note that the dependence of the sample complexity on $d$ and $p$ in Corollary 1 matches that of the lower bound in Theorem 3. Thus, our sample complexity bound in Corollary 1 is optimal with respect to $d$ and $p$.
Remark 3 (Effect of Degree of Separation).
@@
As the parameter quantifying the degree of separation $\gamma$ decreases, the accuracy lower bound in Theorem 2 also decreases and the sample complexity required to outperform the noisy label source in Corollary 1 increases. This is consistent with our intuition that a classification task should become harder as the degree of separation reduces; we also saw this in Figure 1.
@@ 4.2 Retraining
We first label the unlabeled set $\mathcal{X}_{U}:=\{{\mathbf{x}}^{\prime}_{1},\ldots,{\mathbf{x}}^{\prime}_{n}\}$ using the learned classifier $\widehat{\bm{\theta}}_{0}$. Let us denote the predicted label for ${\mathbf{x}}^{\prime}_{j}$ with $\widehat{\bm{\theta}}_{0}$ by $\widetilde{\mathrm{y}}_{j}^{\prime}=\textup{sign}\big{(}\langle{\mathbf{x}}^{\prime}_{j},\widehat{\bm{\theta}}_{0}\rangle\big{)}$. Also, let $\mathcal{T}_{2}:=\{({\mathbf{x}}^{\prime}_{j},\widetilde{\mathrm{y}}_{j}^{\prime})\}_{j\in[n]}$; we use $\mathcal{T}_{2}$ for retraining. Just like Section 4.1 (using the squared loss), our separator of interest with $\mathcal{T}_{2}$ is:
$\widehat{\bm{\theta}}_{1}=\frac{1}{\sigma^{2}}\Bigg{(}\frac{1}{n}\sum_{j=1}^{n}\widetilde{\mathrm{y}}_{j}^{\prime}{\mathbf{x}}_{j}^{\prime}\Bigg{)}.$ (3)
Theorem 4 (Retraining: Probability of Predicted Label = True Label).
Consider a sample ${\mathbf{x}}\notin\mathcal{X}_{L}\cup\mathcal{X}_{U}$. Let $\beta({\mathbf{x}})$ denote the angle between ${\mathbf{x}}$ and $\bm{\theta}^{\ast}$. Also, let $p_{2}:=\exp\Big{(}-\frac{9(d-1)}{40}\Big{)}+\exp\Big{(}-\frac{n\mu^{2}\gamma^{2}}{12d}(1-2p)^{2}\Big{)}$. As long as $|\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}|\geq\frac{16p_{2}}{\mu}$, our predicted label for ${\mathbf{x}}$, viz., $\textup{sign}\big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{1}\rangle\big{)}$ is equal to its ground truth label, viz., $\textup{sign}\big{(}\langle{\mathbf{x}},\bm{\theta}^{\ast}\rangle\big{)}$ with a probability (over the randomness of $\mathcal{T}$ and $\mathcal{T}_{2}$) of at least $\alpha_{1}({\mathbf{x}}):=1-\exp\Big{(}-\frac{n\mu^{2}}{4}(1-2p_{2})\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Big{)}$.
The proof of Theorem 4 is in Appendix E; it involves non-trivial moment generating function (MGF) computations. The analysis is especially challenging because as we discussed after Theorem 1, the classifier learned with vanilla training, i.e., $\widehat{\bm{\theta}}_{0}$, is a non-uniform noisy label provider. Also, just like the classifier $\widehat{\bm{\theta}}_{0}$, the classifier $\widehat{\bm{\theta}}_{1}$ learned with retraining is more likely to be wrong on the samples that are less aligned with the ground truth separator $\bm{\theta}^{\ast}$.
We will now provide a lower bound on the population accuracy of the classifier $\widehat{\bm{\theta}}_{1}$; we denote this by $\text{acc}(\widehat{\bm{\theta}}_{1})$ and it is defined as:
$\text{acc}(\widehat{\bm{\theta}}_{1}):=\mathbb{P}_{{\mathbf{x}}\sim\mathcal{D},\mathcal{T},\mathcal{T}_{2}}\Big{(}\textup{sign}\big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{1}\rangle\big{)}=\textup{sign}\big{(}\langle{\mathbf{x}},\bm{\theta}^{\ast}\rangle\big{)}\Big{)}.$
Theorem 5 (Retraining: Population Accuracy).
Suppose $p>(1+\Theta(1))\exp(-\Theta({d}))$ and $\gamma\geq\Theta\Big{(}\frac{\sqrt{d}\exp(-\Theta(d))}{\mu}\Big{)}$. Let $n\geq\Theta\Big{(}\frac{1}{(1-2p)^{2}}\Big{(}\frac{d}{\mu^{2}\gamma^{2}}\Big{)}\max\Big{(}\log\Big{(}\frac{d}{\mu^{2}\gamma^{2}}\Big{)},\log\frac{1}{p}\Big{)}\Big{)}$. Then:
@@
$\textup{acc}(\widehat{\bm{\theta}}_{1})\geq 1-\exp\big{(}-\Theta(d)\big{)}-\exp\Bigg{(}-\frac{n{\mu^{2}\gamma^{2}}(1-2p_{2})}{\Theta(d)}\Bigg{)},$
with $p_{2}=\exp\big{(}-\Theta(d)\big{)}+\exp\Big{(}-\frac{n\mu^{2}\gamma^{2}(1-2p)^{2}}{\Theta(d)}\Big{)}\leq\min\Big{(}p,\Theta\big{(}\frac{\mu\gamma}{\sqrt{d}}\big{)}\Big{)}$.
The proof of Theorem 5 is in Appendix F. Note that the sample complexity required for Theorem 5 to hold is more than the lower bound in Theorem 3 (specifically, there is an extra $\log\big{(}\frac{d}{\mu^{2}\gamma^{2}}\big{)}$ term in Theorem 5). So Theorem 3 is never violated by Theorem 5.
We will now compare the accuracies of vanilla training and retraining.
Remark 4 (When Does Retraining Improve Accuracy?).
Suppose we are in the non-trivial noise regime where $p$ is a constant (e.g., $\frac{2}{5}$). Moreover, suppose $\gamma=\Theta(\mu)$. Then, as per Theorem 5, when $n\geq\Theta\Big{(}\frac{1}{(1-2p)^{2}}\big{(}\frac{d}{\mu^{4}}\big{)}\log\big{(}\frac{d}{\mu^{4}}\big{)}\Big{)}$, [^6] This follows because $\gamma=\Theta(\mu)$ and $p$ is a constant. the accuracy of retraining is $\geq 1-\exp\big{(}-\Theta(d)\big{)}-\exp\Big{(}-\frac{n\mu^{4}}{\Theta(d)}\Big{)}$. [^7] This holds because $\gamma=\Theta(\mu)$ and $p_{2}\leq\Theta\Big{(}\frac{\mu^{2}}{\sqrt{d}}\Big{)}$, since $p$ is a constant. This can be further lower-bounded by:
$\geq 1-\Theta(1)\exp\Bigg{(}-\min\Bigg{(}\Theta(d),\frac{n\mu^{4}}{\Theta(d)}\Bigg{)}\Bigg{)}.$ (4)
In comparison, the accuracy of vanilla training as per Theorem 2 is:
$\leq 1-\Theta(1)\exp\Bigg{(}-\frac{n\mu^{4}(1-2p)^{2}}{\Theta(d)}\Bigg{)}.$ (5)
From eq. 4 and eq. 5 and as per the lower bound on $n$ stated before, when
$n\in\Bigg{(}\Theta\Bigg{(}\frac{(d/\mu^{4})\log\big{(}{{d}/{\mu^{4}}}\big{)}}{(1-2p)^{2}}\Bigg{)},\Theta\Bigg{(}\frac{(d^{2}/\mu^{4})}{(1-2p)^{2}}\Bigg{)}\Bigg{)}$ (6)
@@
and $p$ is sufficiently close to $\frac{1}{2}$, the lower bound on the accuracy of retraining (in eq. 4) is greater than the upper bound on the accuracy of vanilla training (in eq. 5). This improvement becomes increasingly significant as $p$ approaches $\frac{1}{2}$.
Regarding $n$ in Remark 4. We expect that the minimum value of $n$ needed for retraining to be beneficial should be at least the sample complexity needed for the classifier obtained with vanilla training to outperform the noisy label source used to train it. We computed the latter in Corollary 1; compared to this value, the minimum value of $n$ required for Remark 4 to hold (in eq. 6) has an extra mild factor of $\log\big{(}{\frac{d}{\mu^{4}}}\big{)}$. [^8] Recall that this is when $\gamma=\Theta(\mu)$ and when $p$ is a constant. On the other hand, the upper bound on $n$ in eq. 6 of Remark 4 could be an artifact of our analysis; investigating this is left for future work. [^9] More specifically, this upper bound on $n$ arises as a result of the $\exp\big{(}-\Theta(d)\big{)}$ term in the accuracy bound of retraining. If this can be removed, then the upper bound will not exist.
Remark 5 (Effect of Degree of Separation).
Similar to vanilla training (Remark 3), as the parameter quantifying the degree of separation $\gamma$ decreases, the accuracy lower bound of retraining in Theorem 5 also decreases [^10] Note that $p_{2}$ also depends on $\gamma$ and it is a decreasing function of $\gamma$. and the sample complexity required for this bound to hold increases. Moreover, in Remark 4, we considered the regime of $\gamma=\Theta(\mu)$, i.e., $\gamma$ is “large” enough relative to the mean $\mu$. This tells us that retraining is more beneficial when the degree of separation is large.

### v2

@@ 4.1 Vanilla Training
We will next define the accuracy of a classifier ${\widehat{\bm{\theta}}}$ as
${\sf acc}({\widehat{\bm{\theta}}}):=\mathbb{P}({\rm sign}(\langle\bm{x},{\widehat{\bm{\theta}}}\rangle)=y)\,,$ (4.4)
where the probability is with respect to the randomness in the training set used to learn ${\widehat{\bm{\theta}}}$ as well as $(\bm{x},y)$. Our next result bounds the accuracy of the classifier ${\widehat{\bm{\theta}}}_{0}$ obtained by vanilla training.
Theorem 4.2 (Vanilla Training: Population Accuracy)
We have
${\sf acc}({\widehat{\bm{\theta}}}_{0})$ $\leq 1-\frac{1}{4\sqrt{2\pi}}\exp\left(-160(1+\sqrt{n}(1-2p))^{2}\frac{\gamma^{4}}{\lambda_{\min}^{2}d}\right)\left(1-e^{-d/16}\right)\,\text{ and }$ (4.5)
${\sf acc}({\widehat{\bm{\theta}}}_{0})$ $\geq 1-\frac{1}{2}\exp\left(-\frac{n(1-2p)^{2}\gamma^{4}}{16\lambda_{\max}^{2}d}\right)-e^{-d/8}-2\exp\left(-\frac{n}{32}(1-2p)^{2}\right)\,,$ (4.6)
for an absolute constant $c>0$.
The proof of Theorem 4.2 is in Appendix C.
Remark 4.3 (Tightness of Accuracy Bounds)
@@
When $n\lesssim\frac{d^{2}}{(1-2p)^{2}\gamma^{4}}$, $\gamma\lesssim d^{1/4}$ and $\frac{\lambda_{\max}}{\lambda_{\min}}\lesssim 1$, then the lower and upper bounds for $\textup{acc}(\widehat{\bm{\theta}}_{0})$ in Theorem 4.2 match (up to constant factors).
Based on Theorem 4.2, we have the following corollary.
Corollary 4.4 (Vanilla Training: Sample Complexity)
@@
For any $\delta>2e^{-d/8}+4\exp\left(-\frac{n}{32}(1-2p)^{2}\right)$, having sample size $n\geq 8\lambda_{\max}^{2}\frac{\log{1}/{\delta}}{(1-2p)^{2}}\frac{d}{\gamma^{4}}$ ensures that $\textup{acc}(\widehat{\bm{\theta}}_{0})>1-\delta$. In particular, when the label flipping probability satisfies $p>2e^{-d/8}+4\exp\left(-\frac{n}{32}(1-2p)^{2}\right)$, then $n\geq 8\lambda_{\max}^{2}\frac{\log{1}/{p}}{(1-2p)^{2}}\frac{d}{\gamma^{4}}$ ensures that $\textup{acc}(\widehat{\bm{\theta}}_{0})>1-p$, i.e., our learned classifier $\widehat{\bm{\theta}}_{0}$ has better accuracy than the source providing noisy labels (used to learn $\widehat{\bm{\theta}}_{0}$).
Remark 4.5 (Effect of Degree of Separation)
@@
As the parameter quantifying the degree of separation $\gamma$ decreases, the accuracy bound in Theorem 4.2 also decreases and the sample complexity required to outperform the noisy label source in Corollary 4.4 increases. This is consistent with our intuition that a classification task should become harder as the degree of separation reduces; we also saw this in Figure 1.
We conclude this subsection by deriving an information-theoretic lower bound on the sample complexity of any classifier to argue that ${\widehat{\bm{\theta}}}_{0}$ attains the optimal sample complexity with respect to $d$ and $p$.
Theorem 4.6 (Information-Theoretic Lower Bound on Sample Complexity)
With a slight generalization of notation, let ${\sf acc}({\widehat{\bm{\theta}}};{\bm{{\mu}}})$ denote the accuracy of the classifier ${\widehat{\bm{\theta}}}$, when the ground truth separator is ${\bm{{\mu}}}$. We also consider the case of $\bm{\Sigma}=\mathcal{P}^{\perp}_{\bm{\mu}}$, viz., the projection matrix onto the space orthogonal to ${\bm{{\mu}}}$. For any classifier ${\widehat{\bm{\theta}}}$ learned from $\mathcal{T}:=\{(\bm{x}_{j},\widehat{y}_{j})\}_{j\in[n]}$, in order to achieve $\inf_{{\bm{{\mu}}}\in\mathbb{S}^{d-1}}{\sf acc}({\widehat{\bm{\theta}}};{\bm{{\mu}}})\geq 1-\delta$, the condition $n=\Omega\left(\frac{(1-\delta)}{(1-2p)^{2}}d\right)$ is necessary in our problem setting.
It is worth mentioning that there is a similar lower bound in Gentile and Helmbold, (1998) for a different classification setting. In contrast, Theorem 4.6 is tailored to our setting and moreover, the proof technique is also different and interesting in its own right. Specifically, for the proof of Theorem 4.6, we follow a standard technique in proving minimax lower bounds which is to reduce the problem of interest to an appropriate multi-way hypothesis testing problem; this is accompanied by the application of the conditional version of Fano’s inequality and some ideas from high-dimensional geometry. This proof is in Appendix D.
Remark 4.7 (Minimax Optimality of Sample Complexity)
Note that the dependence of the sample complexity on $d$ and $p$ in Corollary 4.4 matches that of the lower bound in Theorem 4.6. Thus, our sample complexity bound in Corollary 4.4 is optimal with respect to $d$ and $p$.
@@ 4.2 Retraining
We first label the training set using ${\widehat{\bm{\theta}}}_{0}$. Denote by $\widetilde{y}_{i}={\rm sign}(\langle\bm{x}_{i},{\widehat{\bm{\theta}}}_{0}\rangle)$ the predicted label for sample $\bm{x}_{i}$. We then retrain the model using these predicted labels. Our learned classifier here is similar to the one in Section 4.1, except that the observed labels are replaced by the predicted labels. Specifically, our retraining classifier is the following:
${\widehat{\bm{\theta}}}_{1}=\frac{1}{n}\sum_{i}\widetilde{y}_{i}\bm{x}_{i}\,.$ (4.7)
Our retraining classifier’s predicted label for a sample $\bm{x}\in\mathbb{R}^{d}$ is ${\rm sign}(\langle\bm{x},\widehat{\bm{\theta}}_{1}\rangle)$. Our next result bounds the probability of ${\widehat{\bm{\theta}}}_{1}$ correctly classifying a given fixed test point $\bm{x}$.
Theorem 4.8 (Retraining)
Suppose that $\frac{n}{d}>\frac{4\lambda_{\max}}{\gamma^{2}(1-2p)}$ and $nd>\frac{\gamma^{4}}{\lambda_{\max}^{2}}$. Consider $\bm{x}\notin\mathcal{T}$ and let $y$ be its true label. We have
$\mathbb{P}({\rm sign}(\langle\bm{x},{\widehat{\bm{\theta}}}_{1}\rangle)=y)\geq\alpha_{1}(\bm{x})\,,$
where
$\alpha_{1}(\bm{x}):=1$ $-2\exp\left(-\frac{n(1-2q^{\prime})^{2}\langle\bm{x},{\bm{{\mu}}}\rangle^{2}}{64(\langle\bm{x},{\bm{{\mu}}}\rangle^{2}+\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2})}\right)$
$-4\exp\left(-\frac{n}{32}(1-2p)^{2}\right)-\frac{n}{2}\left(\exp\left(-\frac{\gamma^{4}}{8\lambda_{\max}^{2}}(1-2p^{\prime})\frac{n}{d}\right)+e^{-d/16}\right)e^{d/n}\,,$
with
$q^{\prime}=\exp\left(-\frac{n(1-2p)\gamma^{2}}{40\lambda_{\max}}\right)\text{ and }p^{\prime}=\left(1+\frac{3\gamma^{4}}{8\lambda_{\max}^{2}nd}\right)p\,.$
The proof of Theorem 4.8 is in Appendix E.
Technical Challenges. The proof of Theorem 4.8 is especially challenging because as we discussed after Theorem 4.1, the classifier learned with vanilla training, i.e., ${\widehat{\bm{\theta}}}_{0}$, is a non-uniform noisy label provider. In addition, ${\widehat{\bm{\theta}}}_{0}$ depends on all the samples in the training set and hence each predicted label $\widetilde{y}_{i}$ is also dependent on the entire training set. This dependence of the predicted labels on all the data points makes the analysis even more challenging because standard arguments with independence are not applicable here. In contrast, the analysis of ${\widehat{\bm{\theta}}}_{0}$ was technically simpler as it was based on noisy labels $\widehat{y}_{i}$’s that were independent across samples. The high-level proof idea for Theorem 4.8 is that for each sample $\bm{x}_{\ell}=y_{\ell}(1+u_{\ell}){\bm{{\mu}}}+\bm{\Sigma}^{1/2}{\bm{z}}_{\ell}$ in the training set, we carefully curate a dummy label $\widetilde{y}^{\prime}_{\ell}$ which does not depend on ${\bm{z}}_{k}$ for $k\neq\ell$, but depends on all $\{u_{k}\}_{k\in[n]}$; see (E.2) in Appendix E. We show that with high probability the labels predicted by the vanilla estimator ${\widehat{\bm{\theta}}}_{0}$ on the training samples, i.e. ${\rm sign}(\langle\bm{x}_{\ell},{\widehat{\bm{\theta}}}_{0}\rangle)$ are the same as the dummy labels $\widetilde{y}^{\prime}_{\ell}$. Next, under this high probability event, we use the dummy labels instead, which are more amenable to analysis as they are independent of $\{{\bm{z}}_{k}\}_{k\neq\ell}$. At the beginning of Appendix E, we provide an outline of these steps in more detail.
It is worth noting that the lower bound derived on $\alpha_{1}(\bm{x})$ is increasing in $\langle\bm{x},{\bm{{\mu}}}\rangle$. So, just like the vanilla training classifier ${\widehat{\bm{\theta}}}_{0}$, the classifier ${\widehat{\bm{\theta}}}_{1}$ learned with retraining is more likely to be wrong on the samples that are less aligned with the ground truth separator ${\bm{{\mu}}}$. We will now provide a lower bound on the accuracy (defined in (4.4)) of the classifier ${\widehat{\bm{\theta}}}_{1}$.
Theorem 4.9 (Retraining: Population Accuracy)
Under the setting of Theorem 4.8 we have
@@
${\sf acc}({\widehat{\bm{\theta}}}_{1})\geq 1$ $-2\exp\left(-\frac{n(1-2q^{\prime})^{2}\gamma^{4}}{64(\gamma^{4}+2\lambda_{\max}^{2}d)}\right)-2e^{-d/8}$
$-4\exp\left(-\frac{n}{32}(1-2p)^{2}\right)-\frac{n}{2}\left(\exp\left(-\frac{\gamma^{4}}{8\lambda_{\max}^{2}}(1-2p^{\prime})\frac{n}{d}\right)+e^{-d/16}\right)e^{d/n}\,.$ (4.8)
The proof of Theorem 4.9 is in Appendix F. As expected, the lower bound (4.8) is increasing in $\gamma$, because as the degree of separation increases, the classification task becomes easier.
Accuracy of Retraining vs. Vanilla Training. We will now compare the accuracy of vanilla training model ${\widehat{\bm{\theta}}}_{0}$ with that of the retrained model ${\widehat{\bm{\theta}}}_{1}$, identifying regimes where retraining improves the model accuracy. Before proceeding further, it is useful to note that the dominant term in $1-{\sf acc}({\widehat{\bm{\theta}}}_{1})$ is the last term in (4.8) which has $(1-2p^{\prime})$ inside the exponent and $p^{\prime}$ can become arbitrarily close to $p$ as $n$ or $d$ grows. In contrast, the upper bound on ${\sf acc}({\widehat{\bm{\theta}}}_{0})$ in (4.5) has $(1-2p)^{2}$ inside the exponent. This is the main observation which indicates that retraining can improve model accuracy, and this improvement becomes increasingly significant as $p$ approaches $\frac{1}{2}$. We will now formalize this observation by providing sufficient conditions for this improvement to happen.
Remark 4.10 (When Does Retraining Improve Accuracy?)
We consider an asymptotic regime where $n,d\to\infty$. Also, suppose that $\lambda_{\min}$ and $\lambda_{\max}$ remain bounded from above and below by positive constants. We would like to characterize a regime where $1-{\sf acc}({\widehat{\bm{\theta}}}_{1})\leq 1-{\sf acc}({\widehat{\bm{\theta}}}_{0})$, i.e., under what conditions the misclassification error of ${\widehat{\bm{\theta}}}_{1}$ is smaller than that of ${\widehat{\bm{\theta}}}_{0}$. By comparing the bounds in Theorems 4.2 and 4.9, we obtain the following sufficient conditions. There exists an absolute constant $c_{0}$ such that if $p\in(\frac{1}{2}-c_{0},\frac{1}{2})$ and
$\frac{d}{\gamma^{2}(1-2p)}\max\left\{1,\frac{\log n}{\gamma^{2}}\right\}\lesssim n\lesssim\frac{d^{2}}{\gamma^{4}(1-2p)^{2}}\,,$ (4.9)
@@
then the lower bound on the accuracy of retraining (in eq. 4.8) is greater than the upper bound on the accuracy of vanilla training (in eq. 4.5).
Note that as the degree of separation $\gamma$ decreases, the range of $n$ in (4.9) becomes larger but also shifts up, indicating we need more samples (keeping other parameters fixed) to obtain an improvement with retraining.